OpenStudy (anonymous):

The point (2,0) lies on the graph of the function y=2x^2-8x+6

OpenStudy (anonymous):

<3 you forever

OpenStudy (anonymous):

if the question is to examine whether the point is on the line, then you would plug in x and see if you get y

OpenStudy (triciaal):

false when x = 2 y = 2(2)^2-8(2) + 6 = 8 - 16 + 6 = -2 to factor 2x^2 -8x + 6 = 0 (2x - 2)(x - 3)= 0 x = 1, 3 on the graph would be (1, 0) (2, -2) (3, 0)

OpenStudy (anonymous):

my bad. thought it was 4x